# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

### v2

@@ 7.1 Comparison with Previous Approaches
In this subsection, we discuss the differences between the proposed SUSA architecture and previous approaches incorporating depth or textual environmental information. For depth-based methods, such as BEVBert [1] and GridMM [58], depth information is primarily leveraged to map the positions of the corresponding RGB images within the environment, facilitating the construction of local or global grid-based representations. In contrast, our approach directly incorporates depth images as inputs, providing the agent with a clearer perception of the environmental layout. On the other hand, CONSOLE [30] leverages ChatGPT https://openai.com/blog/chatgpt to generate landmark cooccurrence priors, which are then aligned with the agent’s visual observations, while KERM [26] describes the attributes and relationships between objects as textual knowledge. Despite the varied forms of incorporating textual semantics, these methods typically compact textual semantics with corresponding visual views before feeding them to the agent. In contrast, our approach treats the generated textual semantics as a separate environmental representation, enabling the agent to better ground environments with textual instructions, thereby significantly improving both navigation success and object recognition accuracy. In summary, compared to previous approaches, we improve spatial-semantic perception for agent by equipping agent with more alternative environmental representations.
Figure 7: The weights for hybrid representations in the HRF module change during training (see Eq. 6).
@@
We quantitatively illustrate the weights (see Eq. 6) assigned to each branch of the HMF module during training (in Fig. 7) and inference (in Tab. 9). These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation. As shown in Fig. 7, initially, all four representations have approximately equal weights of 0.25 during training, reflecting random initialization and balanced contributions. Over the course of training, the global RGB representation weight $\beta_{4}$ increases, while those of the textual semantic $\beta_{1}$ and depth $\beta_{3}$ modalities decrease. This trend arises because textual semantics and depth provide sparse yet precise information, which is particularly beneficial in the early stages of training. Despite the increasing dominance of the RGB modality, the agent’s reliance on textual semantics and depth remains above 20%, underscoring their essential role in navigation. Tab. 9 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
@@
Table 9: Variation in fusion weights (in Eq. 6) of hybrid representations during inference on R2R. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRF module, respectively.
@@ 7.3 Quality of Generated Salient Landmarks
The quality of the extracted salient landmarks is crucial for the agent to accurately understand the textual semantics of the environment. In Fig. 8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R [3] and REVERIE [40] to evaluate the quality of the generated salient landmarks. While the types and frequencies of these generated landmarks is different from those in the datasets, they can contain and convey most similar sematic (similarities surpass 0.5). Notably, the plug-and-play CLIP text encoder excels in understanding image-related descriptions, while BERT is more adept at comprehending natural language sentences. Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks, instead of BERT.
Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRF module on R2R validation unseen split.
@@ 7.4 Analysis of the Contrastive Loss Weight
Our ablation studies indicate that the contrastive learning loss weight, denoted as $\lambda_{2}$ in Eq. 11, significantly affects model navigation performance. As shown in Fig. 9, we observe that, within a certain range, increasing the contrastive loss weight improves navigation performance. However, when the contrastive loss weight is too high, performance (SR/OSR) declines, thus overemphasizing contrastive learning loss may lead the agent to prioritize modality alignment over core sequential navigation tasks.
ID DSP TSU HRF SR SPL RGSPL
1 46.98/52.51 33.73/36.06 23.03/22.06
2 52.00/52.53 39.21/39.53 26.54/24.97
3 53.17/53.89 38.31/40.89 26.73/26.43
4 55.07/51.97 39.50/37.87 26.70/25.20
5 51.75/54.39 38.86/41.54 26.56/27.31
Table 10: Ablation study on different components of SUSA on the REVERIE (validation/test) unseen split.
7.5 Overall Design on REVERIE
We further supplemented the performance of each key component of SUSA on the REVERIE task. To ensure fairness, we adopt the same hyperparameter setting ($\lambda_{2}$=0.8) as used in R2R, though this may be suboptimal for REVERIE. As shown in Tab. 10, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL). The TSU module enhances the agent’s understanding of textual semantics, resulting in higher success rates in both navigation (SR) and object recognition (RGS). An intriguing phenomenon, as shown in Tab. 10, is observed on the REVERIE dataset: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates. In contrast, the addition of the HRF module (#5) yields opposite results. Overall, the introduction of the HRF module leads to superior navigation performance, both in terms of navigation (SR/SPL) and object grounding (RGSPL).
@@ 7.5 Overall Design on REVERIE
As Fig. 10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e., SR/SPL, which are more important metrics) of different VLN models, including DUET [5] , DSRG [53], HAMT [4] and our SUSA. To ensure a fair comparison across all methods, we performed single-step forward inference with a batch size of 8, an instruction length of 44, and 6 exploration map nodes. Compared to the baseline DUET, our approach introduces spatial and semantic environmental information, which inevitably increasing computational cost. However, despite this modest cost, the SUSA architecture notably enhances agent navigation performance.
@@ 9 Limitations and Future Work
As previously mentioned, when an agent encounters the same landmark at different navigable points, it may struggle to make confident navigation decisions, leading to failures. To address this limitation, 1) our future research will explore adjusting the confidence levels associated with different environmental representations to help the agent make more accurate and reliable predictions. 2) Beyond landmarks, we will construct direction-related textual semantics (e.g., “past,” “turn left”) to better model the spatial relationships between actions in instructions and the environmental layout or historical exploration maps. We also intend to investigate how well our approach generalizes to other navigation tasks. Additionally, given the complexity of VLN tasks, current mainstream VLN frameworks are often intricate and cumbersome, with a wide variety of evaluation metrics that are difficult to balance. Ongoing research in the future is expected to develop more lightweight VLN model architectures and comprehensive evaluation metrics.

### v4

@@ 7.1 Comparison with Previous Approaches
We demonstrate the effectiveness of the proposed method through comparisons with related approaches. As shown in  Tab. 9, experiments reveal that directly grounding depth information—whether using depth panoramas (DP) or depth exploration maps (DEM) as input—outperforms prior depth-based methods [1, 56, 53]. Notably, using depth exploration maps leads to better navigation performance than depth panoramas. Similarly, directly aligning textual semantics with instructions via textual panoramas (TP) surpasses methods [15, 31, 34] that integrate textual semantics into environmental representations. This underscores the superiority of independently grounding each modality, improving the interpretability of each modality’s contribution.
Figure 7: The weights for hybrid representations in the HRA module change during training (see Eq. 6).
@@
We quantitatively illustrate the weights (see Eq. 6) assigned to each branch of the HRA module during training (in Fig. 7) and inference (in Tab. 10). These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation. As shown in Fig. 7, initially, all four representations have approximately equal weights of 0.25 during training, reflecting random initialization and balanced contributions. Over the course of training, the global RGB representation weight $\beta_{4}$ increases, while those of the textual semantic $\beta_{1}$ and depth $\beta_{3}$ modalities decrease. This trend arises because textual semantics and depth provide sparse yet precise information, which is particularly beneficial in the early stages of training. Despite the increasing dominance of the RGB modality, the agent’s reliance on textual semantics and depth remains above 20%, underscoring their essential role in navigation. Tab. 10 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
@@
Table 10: Variation in fusion weights (in Eq. 6) of hybrid representations during inference on R2R. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRA module, respectively.
@@ 7.3 Quality of Generated Salient Landmarks
The quality of the extracted salient landmarks is crucial for the agent to accurately understand the textual semantics of the environment. In Fig. 8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R [2] and REVERIE [41] to evaluate the quality of the generated salient landmarks. While the types and frequencies of these generated landmarks is different from those in the datasets, they can contain and convey most similar sematic (similarities surpass 0.5). Notably, the plug-and-play CLIP text encoder excels in understanding image-related descriptions, while BERT is more adept at comprehending natural language sentences. Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks.
Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRA module on R2R validation unseen split.
@@ 7.4 Analysis of the Contrastive Loss Weight
Our ablation studies indicate that the contrastive learning loss weight, denoted as $\lambda_{2}$ in Eq. 11, affects model navigation performance. As shown in Fig. 9, agent achieves the optimal navigation performance when the $\lambda_{2}$ is set to 0.8. we further observe that, within a certain range, increasing the contrastive loss weight improves navigation performance. However, when the contrastive loss weight is too high, performance (SR/OSR) declines, thus overemphasizing contrastive learning loss may lead the agent to prioritize modality alignment over core sequential navigation tasks.
[…]
1 13.41 3.30 79.86 71.43 60.63
2 13.85 3.08 80.42 72.80 62.71
3 13.78 3.06 81.65 72.92 63.16
4 12.18 3.06 79.99 73.05 64.85
5 14.58 3.08 82.08 73.95 62.84
7.5 Overall Design on R2R
We further supplemented the performance of each key component of SUSA on the R2R task. As shown in Tab. 6, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL). The TSU module is better at understanding textual semantics, resulting in higher success rates. After integrating the DSP and TSU modules, the agent’s performance improved significantly (e.g., SR=73.05, SPL=64.85). When the HRA module is further introduced, the navigation success rate on R2R increases to 73.95, indicating that contrastive learning further enhances instruction-environment alignment. However, its SPL metric remains suboptimal, suggesting that the contrastive learning strategy in the HRA module may excel at accurately aligning the environment with instructions but lacks the capability to optimize the navigation path length.
@@ 7.5 Overall Design on R2R
As Fig. 10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e., SR/SPL, which are more important metrics) of different VLN models, including DUET [6] , DSRG [50], HAMT [5] and our SUSA. To ensure a fair comparison across all methods, we performed single-step forward inference with a batch size of 8, an instruction length of 44, and 6 exploration map nodes. Compared to the baseline DUET, our approach introduces spatial and semantic environmental information, which inevitably increasing computational cost. However, benefiting from the partial pretraining strategy and a shared panorama encoder for both the depth and RGB environmental information, our SUSA architecture notably enhances agent navigation performance with modest computational resources. The frozen BLIP-2 and CLIP models provide the agent with off-the-shelf textual semantics and are excluded from the training or inference.
@@ 9 Limitations and Future Work
As previously mentioned, when an agent encounters the same landmark at different navigable points, it may struggle to make confident navigation decisions, leading to failures. To address this limitation, 1) our future research will explore adjusting the confidence levels associated with different environmental representations to help the agent make more accurate and reliable predictions. 2) Beyond landmarks, we will construct direction-related textual semantics (e.g., “past,” “turn left”) to better model the spatial relationships between actions in instructions and the environmental layout or historical exploration maps.
In the future, we also intend to explore the generalizability of our approach, which integrates richer representations from multiple perspectives, in embodied tasks such as continuous navigation and embodied question answering. Additionally, given the complexity of VLN tasks, current mainstream VLN frameworks are often intricate and cumbersome, with a wide variety of evaluation metrics that are difficult to balance. Ongoing research in the future is expected to develop more lightweight VLN model architectures and comprehensive evaluation metrics.